A car is traveling at 50 ms-¹from rest cover s distance of 10km in 40, minutes, Calculate the acceleration.


1
Expert's answer
2021-05-23T16:34:18-0400

We can find the acceleration of the car from the kinematic equation:


"d=v_0t+\dfrac{1}{2}at^2."

Since the car starts from rest ("v_0=0"), we can write:


"d=\dfrac{1}{2}at^2,""a=\dfrac{2d}{t^2},""a=\dfrac{2\cdot10000\ m}{(40\cdot60\ s)^2}=3.5\cdot10^{-3}\ \dfrac{m}{s^2}."